OpenStudy (lena772):

Aluminum metal reacts with Iron(III) oxide to produce Iron metal and Aluminum oxide. What mass (in grams) of Aluminum metal will react with 1.50E2 g of Iron(III) oxide?

OpenStudy (lena772):

@hartnn @texaschic101

OpenStudy (lena772):

@Nnesha @sammixboo @UsukiDoll @Jack1

OpenStudy (lena772):

\[Fe _{2}O _{3}+2Al \rightarrow Al _{2} O3+2Fe\]

OpenStudy (aaronq):

so you wrote the formula and balanced it. the next step required is to convert the mass given to moles. after, use a molar ratio - of the moles and coefficients of each species: \(\sf \dfrac{moles~of~A}{A's~coefficient}=\dfrac{moles~of~B}{B's~coefficient}\) solve for moles of Al. Convert moles to mass.

OpenStudy (lena772):

What are the moles of each substance @aaronq ?

OpenStudy (aaronq):

The problem tells you the mass of Fe2O3, convert that to moles. The moles of Al are unknown, so you will be solving for them - they would be your "x"

OpenStudy (lena772):

(1.50E2gFe2O3 * 1 mol) / 159.69 g Fe2O3 = 0.939 mol Fe2O3

OpenStudy (lena772):

(0.939 mol Fe2O3 * 2 mol Al) / 1 mol Fe203 = 1.878 mol Al

OpenStudy (lena772):

1.878 mol Al *26.981g Al / 1 mol Al = 50.67g Al

OpenStudy (lena772):

Would 50.8 be correct considering sig figs?

OpenStudy (aaronq):

50.7 g, no?

OpenStudy (lena772):

sorry. so 50.7 is correct?

OpenStudy (aaronq):

yes, it should be